OpenStudy (anonymous):

what would best describe the pattern 10,13,18 ? Thanks :)

OpenStudy (anonymous):

a-1+n, 4+n, 9+n b-n/2, n/3, n/4 c-19-n, 22-n, 25-n d- 2n, 3n , 4n

OpenStudy (anonymous):

a-1+n, 4+n, 9+n b-n/2, n/3, n/4 c-19-n, 22-n, 25-n d- 2n, 3n , 4n

OpenStudy (anonymous):

Looks like a. 9+ 1 = 10 9+ 4 = 13 9 + 9 = 18.